# Optimal time estimation and the clock uncertainty relation for stochastic processes

### v1

@@ I.2.1 Two-state system
I.2.2 Arbitrary number of states
@@
Appendix J Allan variance
@@
where $\mathsf{\Pi}_{k}=\mathbf{r}_{k}\mathbf{l}^{\sf T}_{k}$ is the projector onto $\mathbf{r}_{k}$, the right eigenvector of $\mathsf{L}$ with eigenvalue $\lambda_{k}$, with $\mathbf{l}_{k}^{\sf T}$ the corresponding left (row) eigenvector [44]. The left and right eigenvectors form a biorthonormal set, such that $\mathbf{l}_{k}^{\sf T}\mathbf{r}_{k^{\prime}}=\delta_{kk^{\prime}}$. The second line of Eq. (J) follows from the fact that $\lambda_{0}=0$, with the steady state $\mathbf{p}^{\rm ss}=\mathbf{r}_{0}$ and the constant row vector $\mathbf{u}^{\sf T}=\mathbf{l}^{\sf T}_{0}$ as the corresponding right and left eigenvectors. We therefore have $\mathbf{u}^{\sf T}\tilde{\mathsf{J}}\mathsf{\Pi}_{0}\tilde{\mathsf{J}}\mathbf{p}^{\rm ss}=(\mathbf{u}^{\sf T}\tilde{\mathsf{J}}\mathbf{p}^{\rm ss})^{2}=J^{2}$ from Eq. (81).
Using Eq. (J) in Eq. (J) and performing the integral, we have

### v2

@@ I.3 Observables linear in the waiting times $\tau_{j}$
K.2.2 Arbitrary number of states
@@
Appendix L Allan variance
@@
where $\mathsf{\Pi}_{k}=\mathbf{r}_{k}\mathbf{l}^{\sf T}_{k}$ is the projector onto $\mathbf{r}_{k}$, the right eigenvector of $\mathsf{L}$ with eigenvalue $\lambda_{k}$, with $\mathbf{l}_{k}^{\sf T}$ the corresponding left (row) eigenvector [46]. The left and right eigenvectors form a biorthonormal set, such that $\mathbf{l}_{k}^{\sf T}\mathbf{r}_{k^{\prime}}=\delta_{kk^{\prime}}$. The second line of Eq. (L) follows from the fact that $\lambda_{0}=0$, with the steady state $\mathbf{p}^{\rm ss}=\mathbf{r}_{0}$ and the constant row vector $\mathbf{u}^{\sf T}=\mathbf{l}^{\sf T}_{0}$ as the corresponding right and left eigenvectors. We therefore have $\mathbf{u}^{\sf T}\tilde{\mathsf{J}}\mathsf{\Pi}_{0}\tilde{\mathsf{J}}\mathbf{p}^{\rm ss}=(\mathbf{u}^{\sf T}\tilde{\mathsf{J}}\mathbf{p}^{\rm ss})^{2}=J^{2}$ from Eq. (136).
Using Eq. (L) in Eq. (L) and performing the integral, we have